OpenStudy (anonymous):

I'm having trouble finding the derivative of this function: f(x) = x ( 1 - 4/ x + 3) *Note: The one is not part of the numerator in the fraction it, should read: -4 / x + 3

OpenStudy (anonymous):

f(x)=x-4x/(x+3) d/dx( f(x) ) = 1-4/(x+3)+4x/(x+3)^2 (using product rule on -4/(x+3) )

OpenStudy (anonymous):

where did the 4x come from?

OpenStudy (anonymous):

sorry, its 4/(x+3)^2

OpenStudy (anonymous):

no x.

OpenStudy (anonymous):

So that part should be x / (x+3) ^ 2?

OpenStudy (anonymous):

no just 4/(x+3)^2

OpenStudy (anonymous):

erivativef(x) = D(x ( 1( - 4/ x + 3))? =x-4x/(x+3)? is this it?

OpenStudy (anonymous):

D[x-4x/(x+3)] =1-D[4x/(x+3)]

OpenStudy (anonymous):

=1-[4(x+3)-4x]/(x+3)^2=1-[12/(x+3)^2]

OpenStudy (anonymous):

Is the original equation x times (-4/(x+3))? If so we make the numerator -4x and denominator x+3. d/dx of a fraction.. we make numerator f(x) and denom g(x).. We do (g(x)f'(x) - f(x)g'(x))/(g(x))^2. We get ((x+3)(-4) - (-4x*1))/((x+3))^2 = -12/((x+3)^2)

OpenStudy (anonymous):

good luck jerjason

OpenStudy (anonymous):

i stick by what i said right at the beginning.I got confused midway through the discussion.But my first answer seems alright.